OpenStudy (anonymous):

find f(x) given that the slope of the tangent line to the graph of f at any point is f'(x): f'(x)=eˣ + x and the graph of f passes through (0,3)

OpenStudy (anonymous):

\[f(x)=e^x+\frac{x^2}{2}+C\] \[f(0)=e^0+\frac{0^2}{2}+C=3\] \[1+C=3\]\[C=2\] \[f(x)=e^x+\frac{x^2}{2}+2\]